Fourty (40) students go to a party wearing red, white and blue. Of these students, 17 wear red, 22 wear white, 25 wear blue. (Students do not necessarily wear only one colour.) Furthermore, 7 wear red and white, 12 wear blue and white, and 9 wear red and blue. Which one of the following alternatives is true? (Hint: first calculate the value of x, the unknown) 1. 5 students wear red only. 8 students wear white and blue, but not red. 3 students wear red and white, but not blue. 2. 2 students wear red only. 11 students wear white and blue, but not red. 6 students wear red and white, but not blue. 3. 2 students wear red only. 8 students wear white and blue, but not red. 3 students wear red and white, but not blue. 4. 5 students wear red only. 11 students wear white and blue, but not red. 6 students wear red and white, but not blue.


1
Expert's answer
2021-05-31T17:42:53-0400

Let

"N(R\cap B\cap W)=x"



Then


"1+x+9-x+x+7-x"

"+12-x+4+x+3+x=40"


"x=4"

"1+x=5"

5 students wear red only.

"12 - 4=8"

8 students wear white and blue, but not red.


"7 - 4 = 3"


1. 5 students wear red only. 8 students wear white and blue, but not red. 3 students wear red and white, but not blue.Β